OpenStudy (anonymous):

Simplify the expression root 6 +2root 6

OpenStudy (saifoo.khan):

\[1\sqrt6 + 2 \sqrt 6 \to (1+2) \sqrt 6 \to 3 \sqrt 6\]

OpenStudy (anonymous):

mostly ,3root6

OpenStudy (anonymous):

Thanks a lot guys :))

OpenStudy (saifoo.khan):

Np! (: